# SustainDiffusion: Optimising the Social and Environmental Sustainability of Stable Diffusion Models

### v1

@@ 5. Results
In the following, we report the results of our RQs. For RQ1 and RQ2, we report in Table 3 the mean and standard deviation of each objective metric for all approaches analysed. In brackets, we report whether the $p-$value of the Wilcoxon test between SustainDiffusion and the baseline approach is lower or greater than our confidence level of 0.008 ($0.05/6$). When the test is significant, we also report the $\hat{A}_{12}$ effect size. In the table, for each objective, the best value is highlighted in bold, while the worst is underlined. The distribution of each objective score is shown in Figure 3. For the trade-off analysis, Table 4 reports the number of Pareto-optimal solutions for each strategy analysed. Finally, we show in Figure 4 the distribution of Hypervolume scores for each strategy in the ten runs. Non-statistically significant differences in hypervolumes are highlighted with a * in Figure 4. For RQ3, we report in Table 5 the $p-$values of the Kruskal-Wallis H test computed between the results returned by SustainDiffusion for each objective in the ten runs. Finally, for RQ4, we report in Table 6 the results of the Win-Tie-Loss evaluation between SustainDiffusion and the baseline approaches.
@@ 5.1. RQ1: Baseline Comparison
SD3 Default 0.64$\pm$0.037 $(<0.008-mid)$ $1$$\pm$00 $(<0.008-large)$ $0.76$$\pm$$0.11$ $(<0.008-large)$ $0.0002$$\pm$$2\times 10^{-06}$ $(<0.008-large)$ $0.0019$$\pm$$2.3\times 10^{-05}$ $(<0.008-large)$ $34$$\pm$$0.13$ $(<0.008-large)$
RS $0.67$$\pm$$0.072$ $(>0.008)$ $0.61$$\pm$$0.34$ $(<0.008-large)$ $0.62$$\pm$$0.21$ $(<0.008-large)$ $0.00054$$\pm$$0.00012$ $(<0.008-large)$ $0.0023$$\pm$$0.00058$ $(<0.008-large)$ $41$$\pm$$9.4$ $(<0.008-large)$
No Prompt Eng. $0.77$$\pm$$0.078$ $(>0.008)$ $0.97$$\pm$$0.082$ $(<0.008-large)$ $0.34$$\pm$$0.084$ $(<0.008-mid)$ $0.00035$$\pm$$5.6\times 10^{-05}$ $(<0.008-large)$ $0.00096$$\pm$$0.00016$ $(>0.008)$ $27$$\pm$$4.3$ $(>0.008)$
Img Q. 0.81$\pm$0.023 $(>0.008)$ $0.7$$\pm$$0.33$ $(<0.008-large)$ $0.38$$\pm$$0.13$ $(>0.008)$ $0.00037$$\pm$$5.5\times 10^{-05}$ $(<0.008-large)$ $0.001$$\pm$$0.00015$ $(>0.008)$ $28$$\pm$$4.2$ $(>0.008)$
Img Q. + Bias $0.68$$\pm$$0.067$ $(>0.008)$ $0.37$$\pm$$0.3$ $(>0.008)$ 0.28$\pm$0.11 $(>0.008)$ $0.00037$$\pm$$0.00015$ $(<0.008-large)$ $0.0011$$\pm$$0.00038$ $(>0.008)$ $30$$\pm$$11$ $(>0.008)$
Img Q. + Energy $0.71$$\pm$$0.092$ $(>0.008)$ $0.52$$\pm$$0.29$ $(<0.008-large)$ $0.53$$\pm$$0.13$ $(<0.008-large)$ 0.00012$\pm$ 5.4 $\times$10-05 $(>0.008)$ 0.0007$\pm$0.00033 $(>0.008)$ 20$\pm$9.1 $(>0.008)$
SustainDiff. $0.69$$\pm$$0.066$ 0.32$\pm$0.29 $0.31$$\pm$$0.14$ $0.00015$$\pm$$3.1\times 10^{-05}$ $0.00094$$\pm$$0.0002$ $25$$\pm$$5.3$
@@
The first two rows in Table 3 report the mean and standard deviation of the single objectives for the SD3 baseline and RS approach, while the last row of Table 3 reports the results achieved by SustainDiffusion. We observe that SustainDiffusion achieves statistically better results than the baseline SD3 model in all considered objectives, with a large effect size in 5 out of 6 objectives. On average, SustainDiffusion reduces the gender bias of the baseline SD3 model by 68% and the ethnic bias by 59%. In general, we observe that the baseline SD3 model provides the highest gender and ethnic bias among all strategies analysed. This result is in line with previous research (Fadahunsi et al., 2025; Bano et al., 2025). Regarding energy, we observe a reduction of 25% in CPU Energy, 50.5% in GPU Energy, and 26.5% in Duration with respect to the default SD3 model. Finally, we also observe a small but statistically significant improvement of 5.3% in Image Quality. SustainDiffusion also provides statistically significantly better results than the RS model, with a large effect size, in 5 out of 6 objectives. The only objective in which the difference is not statistically significant is Image Quality, where the results obtained by the two approaches are comparable. Compared to the RS approach, on average, SustainDiffusion reduces the gender bias by 47.5% and the ethnic bias by 50%. Concerning energy, RS is the approach that provides the highest energy consumption among all the strategies analysed. Compared to RS, SustainDiffusion reduces the CPU Energy by 72%, GPU Energy by 59.13%, and Duration by 39.04%. The first two boxes in Figure 3 show the distribution of objectives from the default SD3 model and RS approach, while the last box shows the distribution of objectives from SustainDiffusion. As expected, RS is the strategy providing the highest variability in objective scores, while the baseline SD3 model has the lowest variability, especially in gender bias and energy metrics.
@@
Figure 4. RQ1-RQ2: Hypervolume scores across ten runs
@@ 5.2. RQ2: Ablation Study
Rows from 3 to 6 in Table 3 report the results obtained from the modified versions of SustainDiffusion. First, we observe that the version of SustainDiffusion that does not apply prompt engineering (No Prompt Eng. in Table 3) is unable to improve gender bias, achieving an average value of 0.97 for this objective. This result confirms the need to perform prompt engineering, in addition to hyperparameter tuning, to reduce bias in the images generated.
@@ 5.4. RQ4: Generalisation
By further investigating the results obtained, we observed that Image Quality was the only objective for which there was no statistically significant difference among all strategies (Wilcoxon test $p-$value $<0.008$). [^4] See our replication package for additional statistics and detailed results (rep, 2025). Therefore, we repeated the same experiment without considering the Image Quality objective. Results are reported in Table 6(b) and show that the number of times in which SustainDiffusion overcomes the baseline approaches in all objectives significantly improves. SustainDiffusion overcomes the default SD3 model in all objectives in 58.9% of the times, the RS model in 19.6% of the times, and the SD3 model with Fair prompt style in 66% of the times.
Answer to RQ4: SustainDiffusion overcomes the baseline strategies 38.5% of the times in all objectives and 47.4% of the times in at least half of the objectives. Without considering Image Quality, SustainDiffusion overcomes the baselines 62.2% of the times in all objectives and 36.3% of the times in at least half of the objectives. Thus, its solutions can be effectively generalised to different input prompts.
@@ 6.1. Bias Mitigation
Results from our empirical evaluation confirm that the default SD3 model is significantly biased when generating images of a software engineer (Fadahunsi et al., 2025; Sami et al., 2023; Bano et al., 2025). The gender bias embedded in this model, in particular, is extremely challenging to mitigate. As shown in the answer to RQ2, performing only hyperparameter tuning may not be sufficient to provide a fairer gender distribution. At the same time, as highlighted in the answer to RQ4, explicitly asking the model to provide a fair gender and ethnicity representation when generating an image is still not effective. With SustainDiffusion, we demonstrated how the combination of hyperparameter tuning and prompt engineering is a viable solution to mitigate the bias of SD models, with an average reduction of 68% in gender bias and 59% in ethnic bias compared to the baseline model.
Given the high variability exposed by the Gender Bias objective and, to some extent, the Ethnic Bias objective, we computed the non-parametric Spearman correlation coefficient (Spearman, 1961) to assess any possible relationship among these objectives and the others optimised by SustainDiffusion. Results are reported in our replication package (rep, 2025) and show no high correlation (all values are $<0.5$) of Gender and Ethnic Bias with any of the other objectives analysed. This result highlights how bias is unrelated to image quality or energy consumption and can be optimised without significantly impacting the other objectives.
@@ 6.2. Energy Efficiency
Our experimental evaluation demonstrated that search-based hyperparameter tuning is significantly effective in reducing the energy consumption of SD models. From RQ1 results of, we observed that the default SD3 model consumes on average 0.0021 kWh of energy, combining both CPU and GPU consumption, to generate a single image, which corresponds to charging a phone up to around 7%, according to the U.S. Greenhouse Gas Equivalence Calculator. [^5] https://www.epa.gov/energy/greenhouse-gas-equivalencies-calculator SustainDiffusion on average reduces the overall energy consumption of SD3 to generate a single image by 48% counting both CPU and GPU energy consumption. Additionally, the Spearman correlation test highlighted a significant positive correlation (all values are $>0.98$) between CPU energy, GPU energy and time consumption, meaning that optimising for one of these metrics positively affects the others in a significant manner. Otherwise, we did not observe any significant correlation (all values are $<0.2$) between the energy metrics and the Image Quality, Gender Bias, and Ethnic Bias objectives. This result highlights how reducing the energy consumption of SD models does not directly imply a reduction in image quality or an increase in bias, aligning with what has been observed in a previous study (Gong et al., 2024).
Finally, we remark that a single run of SustainDiffusion required $\sim 20$ hours of execution time on our machine, while a single run of the SD3 model optimised by SustainDiffusion takes on average $\sim 25$ seconds. Therefore, the time required to perform a complete run of SustainDiffusion can be recovered with $\sim 2880$ queries of an SD3 model optimised by SustainDiffusion, which is far lower than the daily number of SD users. [^6] To the best of our knowledge, the latest statistics available reported an average of 10 million daily users of SD in 2022 (noa, 2022). Given the extensive number of images generated by SD in 2023 (Valyaeva, 2023), it is reasonable to think that the number of daily users of SD is even larger by this time. Additionally, as shown by the answers to RQ3 and RQ4, the optimal configurations returned by SustainDiffusion are generally consistent over multiple runs and can be effectively applied to different input prompts. Thus, in a practical scenario, SustainDiffusion can be executed only once, and the optimal models returned can be effectively applied to different use cases.
@@ 7. Threats to Validity
Internal Validity. The fitness functions used by NSGA2 (image quality, gender bias and ethnic bias) are measured using automated tools, which may be imprecise. While a human-in-the-loop process, where humans evaluate objectives at each iteration of the NSGA2, may be more precise, it could be less efficient or even infeasible given the large number of images generated and evaluated by SustainDiffusion. To mitigate this threat, we relied only on automated approaches that have been effectively adopted in previous work to assess image quality and bias with a high level of accuracy (Berger et al., 2023; Gong et al., 2024; Fadahunsi et al., 2025). Additional metrics and approaches to automatically evaluate image quality and bias could be assessed in future work. Another threat concerns the binary gender and limited ethnicity classification addressed in our study. While we agree that the classifications considered in our study partially reflect reality, we also argue that automatically assessing non-binary genders and multiple ethnicities in artificially generated images could be more challenging and error-prone, as also highlighted by previous studies (Fadahunsi et al., 2025; Sami et al., 2023; Luccioni et al., 2023). The energy measurement poses another threat to the internal validity of our study as measurements may be influenced by other processes running in the background. Although we cannot fully ensure that no other process was running in the background, we made sure to kill all main and heavy processes and followed standard guidelines by including an idle time of one minute between each run of SustainDiffusion.
@@
Construct Validity. Search-based algorithms and text-to-image generation models are stochastic by nature. To mitigate this threat, we performed multiple runs for each algorithm employed in our evaluation and generated multiple images for each input prompt. Finally, we adopted specific statistical tests, post-hoc effect sizes, and Type I correction methods to assess the obtained results.
External Validity. Our study focuses on a specific version of SD and a set of 56 prompts. To mitigate this threat, we used the latest version of SD available at the time of this study. However, the hyperparameters considered in SustainDiffusion are also available in previous versions of SD. Thus, SustainDiffusion can also be successfully applied in previous versions of this model. As for the prompts, we used a dataset that has been shown to expose significant gender and ethnic bias in previous studies (Fadahunsi et al., 2025; Sami et al., 2023). We also argue that there is still a lack of datasets of prompts that can be adopted to benchmark the gender and ethnic bias of text-to-image generation models.
@@ 8. Conclusion and Future Work
Future works concern the adoption and comparison of different search strategies, like NSGA-3 (Deb and Jain, 2014) or Weighted Sum of all objectives. Additionally, future research can investigate the impact of adopting different approaches to measure image quality, assess gender and ethnic bias, and measure energy to compute the fitness objectives. Finally, this study can be extended to multiple text-to-image generation models and input prompts.

### v2

@@ 5. Results
In this section, we present and discuss the results of all research questions addressed in our work.
@@ 5.1. RQ1: Baseline Comparison
[…]
@@
The first two rows in Table 3 report the mean and standard deviation of the single objectives for the SD3 baseline and RS approach, while the last row of Table 3 reports the results achieved by SustainDiffusion. In brackets, we report whether the $p$-value of the Wilcoxon test between SustainDiffusion and the baseline approach is lower or greater than our confidence level of 0.008 ($0.05/6$). When the test is significant, we also report the $\hat{A}_{12}$ effect size. In the table, for each objective, the best value is highlighted in bold, while the worst is underlined.
We observe that SustainDiffusion achieves statistically better results than the baseline SD3 model in all considered objectives, with a large effect size in 5 out of 6 objectives. On average, SustainDiffusion reduces the gender bias of the baseline SD3 model by 68% and the ethnic bias by 59%. In general, we observe that the baseline SD3 model provides the highest gender and ethnic bias among all strategies analysed. This result is in line with previous research (fadahunsi_how_2025; bano2025doessoftwareengineerlook). Regarding energy, we observe a reduction of 25% in CPU Energy, 50.5% in GPU Energy, and 26.5% in Duration with respect to the default SD3 model. Finally, we also observe a small but statistically significant improvement of 5.3% in Image Quality. SustainDiffusion also provides statistically significantly better results than the RS model, with a large effect size, in 5 out of 6 objectives. The only objective in which the difference is not statistically significant is Image Quality, where the results obtained by the two approaches are comparable. Compared to the RS approach, on average, SustainDiffusion reduces the gender bias by 47.5% and the ethnic bias by 50%. Concerning energy, RS is the approach that provides the highest energy consumption among all the strategies analysed. Compared to RS, SustainDiffusion reduces the CPU Energy by 72%, GPU Energy by 59.13%, and Duration by 39.04%.
The first two boxes in Figure 3 show the distribution of objectives from the default SD3 model and RS approach, while the last box shows the distribution of objectives from SustainDiffusion. As expected, RS is the strategy providing the highest variability in objective scores, while the baseline SD3 model has the lowest variability, especially in gender bias and energy metrics.
@@
Figure 4. RQ1-RQ2: Hypervolume scores across ten runs. Non-statistically significant differences in Hypervolumes are highlighted with a *.
@@ 5.2. RQ2: Ablation Study
Rows from 3 to 6 in Table 3 report the results obtained from the modified versions of SustainDiffusion, together with the Wilcoxon test $p$-value and corresponding effect size.
First, we observe that the version of SustainDiffusion that does not apply prompt engineering (No Prompt Eng. in Table 3) is unable to improve gender bias, achieving an average value of 0.97 for this objective. This result confirms the need to perform prompt engineering, in addition to hyperparameter tuning, to reduce bias in the images generated.
@@ 5.4. RQ4: Generalisation
By further investigating the results obtained, we observed that Image Quality was the only objective for which there was no statistically significant difference among all strategies (Wilcoxon test $p$-value <0.008<0.008). [^4] See our replication package for additional statistics and detailed results (repl_package). Therefore, we repeated the same experiment without considering the Image Quality objective. Results are reported in Table 6(b) and show that the number of times in which SustainDiffusion overcomes the baseline approaches in all objectives significantly improves. SustainDiffusion overcomes the default SD3 model in all objectives in 58.9% of the times, the RS model in 19.6% of the times, and the SD3 model with Fair prompt style in 66% of the times.
Answer to RQ4: SustainDiffusion overcomes the baseline strategies 38.5% of the time in all objectives and 47.4% of the time in at least half of the objectives. Without considering Image Quality, SustainDiffusion overcomes the baselines 62.2% of the time in all objectives and 36.3% of the time in at least half of the objectives. Thus, its solutions can be effectively generalised to different input prompts requesting human-like figures in different professions.
@@ 6.1. Bias Mitigation
Results from our empirical evaluation confirm that the default SD3 model is significantly biased when generating images of a software engineer (fadahunsi_how_2025; sami_case_2023; bano2025doessoftwareengineerlook). The gender bias embedded in this model, in particular, is extremely challenging to mitigate. As shown in the answer to RQ2, performing only hyperparameter tuning may not be sufficient to provide a fairer gender distribution. At the same time, as highlighted in the answer to RQ4, explicitly asking the model to provide a fair gender and ethnicity representation when generating an image is still not effective. With SustainDiffusion, we demonstrated how the combination of hyperparameter tuning and prompt engineering is a viable solution to mitigate the bias of SD models, with an average reduction of 68% in gender bias and 59% in ethnic bias compared to the baseline model.
Given the high variability exposed by the Gender Bias objective and, to some extent, the Ethnic Bias objective, we computed the non-parametric Spearman correlation coefficient (spearman1961proof) to assess any possible relationship among these objectives and the others optimised by SustainDiffusion. Results are reported in our replication package (repl_package) and show no high correlation (all values are <0.5<0.5) of Gender and Ethnic Bias with any of the other objectives analysed. This result highlights how bias is unrelated to image quality or energy consumption and can be optimised without significantly impacting the other objectives.
@@ 6.2. Energy Efficiency
Our experimental evaluation demonstrated that search-based hyperparameter tuning is significantly effective in reducing the energy consumption of SD models. From the results of RQ1, we observed that the default SD3 model consumes on average 0.0021 kWh of energy, combining both CPU and GPU consumption, to generate a single image, which corresponds to charging a phone up to around 7%, according to the U.S. Greenhouse Gas Equivalence Calculator. [^5] https://www.epa.gov/energy/greenhouse-gas-equivalencies-calculator SustainDiffusion on average reduces the overall energy consumption of SD3 to generate a single image by 48% counting both CPU and GPU energy consumption. Additionally, the Spearman correlation test highlighted a significant strong positive correlation (all values are >0.98>0.98) between CPU energy, GPU energy and time consumption, meaning that optimising for one of these metrics positively affects the others in a significant manner. Otherwise, we did not observe any strong correlation (all values are <0.2<0.2) between the energy metrics and the Image Quality, Gender Bias, and Ethnic Bias objectives. This result highlights how reducing the energy consumption of SD models does not directly imply a reduction in image quality or an increase in bias, aligning with what has been observed in a previous study (gong2024greenstableyolooptimizinginferencetime).
Finally, we remark that a single run of SustainDiffusion required $\sim 20$ hours of execution time on our machine, with an energy consumption (counted as CPU plus GPU energy) of $\sim 4.071$ kWh. In contrast, a single run of the SD3 model optimised by SustainDiffusion takes on average $\sim 25$ seconds, with an energy consumption of $\sim 0.00101$ kWh. Therefore, the time required to perform a complete run of SustainDiffusion can be recovered with $\sim~2880$ queries of an SD3 model optimised by SustainDiffusion, while the energy consumption can be recovered with $\sim 4030$ prompts. While these numbers may seem high, they are far lower than the daily number of SD users. [^6] To the best of our knowledge, the latest statistics available reported an average of 10 million daily users of SD in 2022 (noauthor_digital_2022). Given the extensive number of images generated by SD in 2023 (noauthor_ai_2023), it is reasonable to think that the number of daily users of SD is even larger by this time. Additionally, as shown by the answers to RQ3 and RQ4, the optimal configurations returned by SustainDiffusion are generally consistent over multiple runs and can be effectively generalised to different input prompts depicting human-like figures in different professions. Thus, in a practical scenario, SustainDiffusion can be executed only once, and the optimal models returned can be effectively applied to different use cases. Finally, we note that all our experiments have been executed on a domestic machine with an entry-level GPU. On the contrary, fine-tuning or compressing an SD model may require additional resources, not only in terms of computational power but also in terms of data and specialised knowledge.
@@ 7. Threats to Validity
The fitness functions used by NSGA2 (image quality, gender bias and ethnic bias) are measured using automated tools, which may be imprecise. Although a human-in-the-loop process, in which humans evaluate the objectives at each iteration of the NSGA2, may be more precise, it could be less efficient or even infeasible given the large number of images generated and evaluated by SustainDiffusion. To mitigate this threat, we relied only on automated approaches that have been effectively adopted in previous work to assess image quality and bias with a high level of accuracy (berger2023stableyolo; gong2024greenstableyolooptimizinginferencetime; fadahunsi_how_2025). Additional metrics and approaches to automatically evaluate image quality and bias could be assessed in future work. Another threat concerns the binary gender and limited ethnicity classification addressed in our study. We acknowledge that the classifications considered in our study may partially reflect reality, however automatically assessing non-binary genders and multiple ethnicities in artificially generated images could be more challenging and error-prone, as also highlighted by previous studies (fadahunsi_how_2025; sami_case_2023; luccioni_stable_2023). The energy measurement poses another threat to the internal validity of our study as measurements may be influenced by other processes running in the background. Although we cannot fully ensure that no other process was running in the background, we made sure to kill all main and heavy processes and followed standard guidelines by including an idle time of one minute between each run of SustainDiffusion.
Construct Validity.
@@
Search-based algorithms and text-to-image generation models are stochastic by nature. To mitigate this threat, we performed multiple runs for each algorithm employed in our evaluation and generated multiple images for each input prompt. Finally, we adopted specific statistical tests, post-hoc effect sizes, and Type I correction methods to assess the obtained results.
External Validity.
Our study focuses on a specific version of SD and a set of 56 prompts. To mitigate this threat, we used the latest version of SD available at the time of this study. However, the hyperparameters considered in SustainDiffusion are also available in previous versions of SD. Thus, SustainDiffusion can also be successfully applied in previous versions of this model. Additionally, SustainDiffusion could also be used with different image-generation models employing the Diffuser architecture. For instance, Flux.1, which is the second most adopted open-source text-to-image generation model after StableDiffusion (flux), includes guidance scale and inference steps as tunable hyperparameters. One adaptation may be the implementation of prompt weighting, which in Flux is managed differently (flux). Nevertheless, we believe such an approach is feasible. As for the prompts, we used a dataset that has been shown to expose significant gender and ethnic bias in previous studies (DALOISIO2026107956; fadahunsi_how_2025; sami_case_2023). While the prompts in this dataset are similar in structure, they allow us to generalise the results to all use cases involving human-like figures in different tasks. A complete analysis of the generalisation of SustainDiffusion under different prompts and datasets is an important area for future research. We provide a replication package and strive to report the details of our study design to allow for reproducibility, replicability and extension of our work (williams2026NIER).
@@ 8. Conclusion and Future Work
Future works concern a more extensive evaluation of SustainDiffusion generability using multiple and diverse prompts, as well as a qualitative evaluation, in terms of quality and fairness, of the images generated by an SD model optimised by SustainDiffusion. Future research can also investigate the adoption and comparison of different search strategies, like NSGA-3 (deb2014evolutionary) or Weighted Sum of all objectives. Finally, future studies can explore the impact of adopting different approaches to measure image quality, assess gender and ethnic bias, and measure energy to compute the fitness objectives.
Acknowledgements.
G. d’Aloisio is partially funded by the European Union – NextGenerationEU through the Italian Ministry of University and Research, Projects PRIN 2022 PNRR “FRINGE: context-aware FaiRness engineerING in complex software systEms” grant n.P2022553SL.